OpenStudy (anonymous):

Two cars leave towns 320 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour less than the other's. If they meet in 2 hours, what is the rate of the slower car? Do not do any rounding.

OpenStudy (jacob902):

Two cars leave towns 750 kilometers apart at the same time and travel toward each other...HELPP!!? Two cars leave towns 750 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour less than the other's. If they meet in 5 hours, what is the rate of the slower car? Do not do any rounding. _______________KILOMETERS PER HOUR

OpenStudy (jacob902):

Two cars leave towns 750 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour less than the other's. If they meet in 5 hours, what is the rate of the slower car? D=RT 750 = [(R - 18) + R]5 750 = (2R - 18)5 750 = 10R - 90 10R = 750 + 90 10R = 840 R = 84 THE SPEED OF THE SLOWER CAR: 66 KILOMETERS PER HOUR

OpenStudy (anonymous):

that was wrong... the correct answer was 71 :/